OpenStudy (anonymous):

What is the slope of the line 6x + 8y = 12?

OpenStudy (johnweldon1993):

You need to turn this equation....into something that looks like y = mx + b *notice the 'y' is by itself on the left side...so how would you get 'y' by itself on the left side in your equation?

OpenStudy (raden):

alternative : if given a straight line ax + by = c, then the slope always m = -a/b

OpenStudy (johnweldon1993):

Yes, or that haha....just always like making it in slope intercept form, easier to read for me..but good point @RadEn

OpenStudy (anonymous):

you subtract??

OpenStudy (anonymous):

im supposed to be in 9th grade cause i skipped a couple grades and im having a really hard time with this

OpenStudy (johnweldon1993):

Actually, just to make it easier on this exact question...use what RadEn has shown. You have an equation ax + bx = c to find the slope.... you do slope = -a --- b So your equation, what is 'a' and what is 'b'? 6x + 8y = 12

OpenStudy (anonymous):

im in a 11 grade class

OpenStudy (anonymous):

x=2-4y/3?

OpenStudy (raden):

6x + 8y = 12 here known a=6 and b=8 m = -a/b = -6/8 = -3/4

OpenStudy (anonymous):

????

OpenStudy (raden):

a is the coef of x, while b is the coef of y m = -a/b = - .../... still confused ?